OpenStudy (anonymous):

A cat is able to land on its feet after a fall,taking the advantage of the principle of conservation of angular momentum.Explain how ?

OpenStudy (nottim):

Well, momentum is p=mv....

OpenStudy (nottim):

Maybe since d=vt, the cat is somehow decreasing the distance or time by falling a special way?

OpenStudy (nottim):

p=m(d/t) or... p=(F/a)(d/t)

OpenStudy (nottim):

if you think f=ma. I'm using super simple physics right now...

OpenStudy (mani_jha):

A cat stretches its body, increasing its moment of inertia. As the moment of inertia, the angular velocity decreases.(L=Iw is constant) So, the cat turns slowly, and can easily turn over and land on its feet. If however the turn is faster, the cat may have a higher chance of falling on its back than on its feet.

OpenStudy (anonymous):

yes wat manijha says is perfectly correct but do remeber that as he says this is possible only on the assumption that there isnt much of a height to which the cat leaps(as gravity creates a torque on the cat as it is an extended body and angular momentum is conserved only when there is no net torque on the system)